OpenStudy (anonymous):

what is the solution to the equation? -4x+5(x-2) =-16+10

OpenStudy (fanduekisses):

First of all, distribute the 5(x-2)

OpenStudy (anonymous):

-4x + 5(x - 2) = -16 +10 -4x + 5x -10 = -6 1x - 10 = -6 +10 +10 1x = 4 /1 /1 x = 4